Functions f , g : R → R are defined, respectively, by f (x) = x2 + 3x + 1, g (x) = 2x – 3,
find   f o f  

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given that,
f(x)=x^2+3x+1,
g(x)=2x–3

fof = f(f(x))

=f(x^2+3x+1)
=(x^2+3x+1)^2+3(x^2+3x+1)+1
=x^4+ 9x^2+ 1+ 6x^3+ 6x + 2x^2+ 3x^2+ 9x + 3 + 1
=x^4+ 6x^3+14x^2+15x + 5


Thanks